### v1

@@ Abstract
Spherical surface parameterization is a fundamental tool in geometry processing and imaging science. For a genus-0 closed surface, many efficient algorithms can map the surface to the sphere; consequently, a broad class of task-driven genus-0 mapping problems can be reduced to constructing a high-quality spherical self-map. However, existing approaches often face a trade-off between satisfying task objectives (e.g., landmark or feature alignment), maintaining bijectivity, and controlling geometric distortion. We introduce the Spherical Beltrami Differential (SBD), a two-chart representation of quasiconformal self-maps of the sphere, and establish its correspondence with spherical homeomorphisms up to conformal automorphisms. Building on the Spectral Beltrami Network (SBN)[xu2025neural], we propose a neural optimization framework BOOST that optimizes two Beltrami fields on hemispherical stereographic charts and enforces global consistency through explicit seam-aware constraints. Experiments on large‑deformation landmark matching and intensity‑based spherical registration demonstrate the effectiveness of our proposed framework. We further apply the method to brain cortical surface registration, aligning sulcal landmarks and jointly matching cortical sulci depth maps, showing improved task fidelity with controlled distortion and robust bijective behavior.
@@
surface mesh, quasiconformal mapping, surface parametrization, deep learning
@@ 1 Introduction
Diffeomorphic mappings are key in computational geometry and medical imaging for surface registration and shape analysis, preserving topology and geometry. Mathematically, a mapping $f:\mathcal{M}\rightarrow\mathcal{N}$ between Riemanian manifolds $\mathcal{M}$ and $\mathcal{N}$ is diffeomorphic if it is a bijection with a smooth inverse. Among all practices, the 3D genus-0 closed surface mapping problem gains attention due to its relevance in the real world, as well as the lack of accurate and generic methods. For example, brain surface registration enables standardized comparisons between individuals for statistical analysis [lui2007landmark]. Over the past decades, numerous works have been proposed to solve this problem. In terms of data form, they can be divided into two main classes, surface mesh [cohen2018spherical, hanocka2019meshcnn, choi2015flash, milano2020primal] and point cloud [lai2025point, zhao2019spherical, zhao2022fast, choi2016spherical].
By the uniformization theorem, the problem of optimizing a mapping from a genus-0 closed surface to another one is equivalent to finding an optimized mapping from the sphere to itself, satisfying certain properties. Established numerical tools [choi2015flash, gu2004genus, crane2013robust], provide conformal spherical parameterizations, ensuring angle preservation. However, a conformal sphere parameterization is usually not desired. For example, some focus on area preserving [lyu2024spherical, choi2022adaptive] and some are interested in finding parameterizations best aligning features [baden2018mobius, guo2023automatic]. Our research objective is to provide a generic and accurate algorithm for this diffeomorphic spherical parameterization problem, formulated mathematically as finding an optimal sphere self-mapping such that
@@
where $E$ is the energy functional by which the solution would satisfy desired properties, and $\mathcal{C}$ is the constraint space in which two factors are indispensable: bijectivity and geometric regularity. For example, in the landmark constrained conformal mapping problem[lui2010optimized, lam2014landmark], to get a bijective and plausible deformation, it is necessary to control the local geometric distortion like angle distortion. Across real-life applications, simultaneous satisfaction of feature alignment, bijectivity, and bounded geometric distortionis crucial but difficult. High-fidelity feature matching tends to introduce severe local stretching while enforcing diffeomorphism and controlled geometric distortion counteracts this tendency. Balancing these terms yields a highly nonconvex landscape with many spurious minima.
Existing methods can be broadly categorized into two groups. The first group optimizes vertex positions directly on the 3D surface. These approaches, which include techniques like Ricci flow [jin2008discrete] and harmonic energy minimization [lai2014folding], are intuitive but suffer from two major drawbacks: the optimization landscape is highly non-convex, and there is no inherent guarantee of bijectivity, sometimes leading to self-intersections. The second group performs optimization on a planar domain[isenburg2001connectivity, kharevych2006discrete, choi2022adaptive, choi2015flash], typically the complex plane $\mathbb{C}$, after an initial stereographic projection. While this simplifies the optimization in some respects, it also faces some issues. For example, boundary constraints are needed because it is a single-chart spherical parameterizations while the unit sphere cannot be covered by a single chart. Additionally, it might induce significant distortion that occurs around the pole of the projection.
A diffeomorphism can be characterized not only by where each vertex moves but also by how the map locally shears and stretches. Considering this, it is natural to explore representing a diffeomorphism by quantities measuring geometric distortion. For a Riemann surface, the Beltrami differential $\mu=\{\mu_{\alpha}\}$ lives in the complex tangent bundle and measures conformal distortion. To overcome limitations of traditional work, we propose representing the self-diffeomorphism of $\mathbb{S}^{2}$ by spherical Beltrami differential, a streamlined variant tailored to the sphere. Given this differential, we recover the map via the Least-Squares Quasiconformal (LSQC) energy [qiu2019computing]. Moreover, numerical implementation of LSQC enjoys properties like existence and uniqueness of minimizer, correspondence with homeomorphisms, invariant by similarity transformation, and independent of resolution of the mesh. However, differentiating through the classical LSQC solver is computationally prohibitive. To overcome this, in the previous work [xu2025neural] we introduce Spectral Beltrami Network (SBN), a deep neural network designed to act as a differentiable surrogate for the LSQC solver.
@@
To solve the diffeomorphic spherical parameterization problem, we propose a novel algorithm, Beltrami Optimization On Spherical Topology (BOOST). BOOST represents a spherical map using two stereographic charts (hemispheres), optimizes the corresponding Beltrami fields using the SBN. The spherical Beltrami differential formalism provides the mathematical object these two fields represent: a single global spherical quasiconformal map once cross-chart compatibility is satisfied. In practice, BOOST enforces this compatibility through seam boundary matching, folding penalties after stitching, and seam smoothness regularization. This two-chart formulation avoids the pole distortions inherent in single-chart spherical parameterizations and enables flexible, distortion-controlled diffeomorphic registration on genus-0 surfaces, with applications to cortical surface alignment and hybrid landmark/feature matching.
@@
The organization of the paper is as follows. In Sec.2, previous related works are reviewed. and contributions are mentioned in Sec.3. In Sec.4, some of the related mathematical concepts are introduced. In Sec.5, we propose the concept spherical Beltrami differential and show its correspondence with spherical quasiconformal mappings, and introduce the model Spectral Beltrami Network. In Sec.6, We establish the neural optimization framework BOOST. In Sec.8, multiple experiments are conducted to show the capability and potential of our work in complex imaging problems. In Sec.9, we analyze the key factors hidden behind the excellence of the model and briefly discuss its limitations as well as potential future work. In Sec.10, a conclusion is given.
@@ 2 Previous works
The spherical parameterization problem of genus-0 closed surfaces has been widely studied, with methods broadly categorized into optimization on the spherical surface and optimization on the parameterization domain.
Gu and Yau in [gu2002computing] proposed a nonlinear gradient descent algorithm to compute global spherical conformal parameterizations for genus-0 surfaces, which is done in the tangent spaces of the sphere, but often converges slowly and cannot guarantee bijectivity. Later, Lai et al.[lai2014folding] accerlerated the algorithm and remove foldings by applying the weighted Laplace–Beltrami eigen-projection. Several flow-based methods are proposed. Jin et al.[jin2008discrete] used Ricci flow to conformally deforms a genus-0 surface to a sphere, but Ricci flow is not strictly convex, which merely settles at a local minimum. Crane et al.[crane2013robust] proposed Willmore flow for triangulated surfaces and applied it to spherical surface fairing. Recent years have seen the emergence of deep learning approaches for spherical surface data. Spherical CNNs [cohen2018spherical, esteves2023scaling] used rotation-equivariant convolutions to analyze spherical signals (e.g., climate models, molecular data), while Spherical U-Net [zhao2019spherical, zhao2022fast] adapts U-Net for tasks like cortical surface parcellation in neuroimaging, employing spherical operations like surface convolution. These methods often treat surfaces as unstructured point sets, neglecting manifold geometry and topological properties, which cannot guarantee bijectivity. For instance, Zhao et al. [zhao2022fast] attempted bijectivity by learning the spherical diffeomorphic deformation field, but discretization errors can result in nonpositive Jacobian determinants, undermining strict bijectivity. Additionally, these approaches are typically task-specific and data-driven, and struggle with large-scale distortions or multiobjective optimization (e.g., balancing conformality and area preservation).
On the other hand, there are also works optimizing spherical self-mappings on the parametrization domain like [lui2012optimization, lui2015splitting, isenburg2001connectivity, kharevych2006discrete, choi2015flash, lyu2024spherical]. Isenburg et al.[isenburg2001connectivity] and Kharevych et al.[kharevych2006discrete] cut meshes into two sections and obtained spherical results using stereographic mapping from plane to sphere. Choi et al. [choi2015flash] based on quasiconformal theory, proposed a north pole–south pole iterative scheme to efficiently compute bijective landmark-aligned spherical harmonic parameterizations for genus-0 closed brain surfaces. Lyu et al. [lyu2024spherical] applied similar ideas to computing spherical density equalizing parameterization for genus-0 surfaces. In the context of deep learning , Guo et al. [guo2023automatic] combined quasi-conformal geometry with CNNs to map triangular meshes with disk topology. They mapped two hemispheres to two new hemispheres and glued them seamlessly. However, these methods often impose fixed boundary constraints, such as prescribing boundaries in outer regions [choi2015flash, lyu2024spherical] or requiring unchanged target domain geometry [guo2023automatic]. These constraints limit flexibility when adapting to complex surface features or tasks that require significant stretching.
In summary, existing methods face limitations in flexibility, theoretical rigor, or generality. Numerical approaches lack adaptability due to fixed boundaries, whereas deep learning methods often ignore manifold structure and lack diffeomorphism guarantees. Our work builds on these insights, particularly improving mesh-based neural approaches to address these gaps effectively.
@@ 3 Contributions
Our work has several key contributions and novel aspects to the field of spherical parameterization, addressing the limitations of existing methods with a principled, free-boundary approach. Specifically, our contributions are as follows:
@@
    BOOST: Beltrami Optimization On Spherical Topology for the large-deformation spherical mapping. We introduce BOOST, a neural optimization framework for spherical/genus-0 surface mapping that splits the sphere into two hemispherical charts, optimizes the associated Beltrami fields in a free-boundary setting using a differentiable quasiconformal solver surrogate, and enforces global consistency via explicit boundary-matching, folding-avoidance, and seam-smoothing objectives. This divide-and-conquer strategy enables robust large deformations without imposing artificial fixed boundary constraints.
@@
    Two-chart spherical Beltrami representation (spherical Beltrami differential) to formalize global consistency. We define the spherical Beltrami differential as a pair of Beltrami fields on two stereographic charts together with the appropriate compatibility across charts, providing a concise representation of spherical self-quasiconformal maps (modulo conformal automorphisms) and a principled lens for controlling local distortion on the sphere.
@@
    Beltrami-space optimization viewpoint for spherical diffeomorphic mapping problems. We formulate spherical mapping tasks as direct optimization over Beltrami fields (rather than vertex coordinates), where the optimized variables encode intrinsic distortion. This leads to a flexible objective design that naturally supports diverse task terms (e.g., landmark alignment, feature-based matching) while maintaining distortion control.
As detailed in Sec.8, this framework successfully handles diverse applications including landmark matching and hybrid registration tasks, demonstrating its use in optimizing spherical self-mappings with bounded distortion and its robustness across challenging spherical mapping scenarios.
@@ 4.1 Quasiconformal Mappings
(2) $\frac{\partial f}{\partial\bar{z}}=\mu(z)\frac{\partial}{\partial z}$
for some measurable complex-valued function $\mu$ with $\|\mu\|_{\infty}<1$, where $\frac{\partial f}{\partial z}=\frac{1}{2}(f_{x}-if_{y})$ and $\frac{\partial f}{\partial\bar{z}}=\frac{1}{2}(f_{x}+if_{y})$. $\mu$ is called the complex dilation or Beltrami coefficient of $f$, which measures the local deviation from a conformal map. Infinitesimally, around a point $p$, $f(p+z)$ might be expressed as an affine map with negligible error $o(|z|)$:
@@
Let $f\colon(x,y)\rightarrow(u,v)$ and $\mu=\rho+i\tau$, and denote by $\mathbf{S}(2)$ the space of all $2\times 2$ symmetric positive definite matrices whose determinant is 1, then
@@
where $Q=(Q_{ij})\colon\Omega\rightarrow\mathbf{S}(2)$. Right multiplying both sides by $D_{f}(z)^{-1}$, we obtain a linear system which is the alternative formulation of the Beltami equation:
@@
and from this we have $\mu=\frac{q_{11}-q_{22}+2iq_{12}}{q_{11}+q_{22}+2\text{sgn}(\det{D_{f}})}$. With the constraint that $Q(z)\in\mathbf{S}(2)$, we observe that $|\mu(z)|<1$ if and only if $\det D_{f}(z)>0$. By inverse function theorem, researchers would through restricting $\|\mu\|_{\infty}<1$, achieve local bijectivity and prevent folding in the triangular mesh [lui2014teichmuller, choi2015flash].

### v2

@@ Abstract
Many genus-0 surface mapping tasks such as landmark alignment, feature matching, and image-driven registration, can be reduced (via an initial spherical conformal map) to optimizing a spherical self-homeomorphism with controlled distortion. However, existing works lack efficient mechanisms to control the geometric distortion of the resulting mapping. To resolve this issue, we formulate this as a Beltrami-space optimization problem, where the angle distortion is encoded explicitly by the Beltrami differential and bijectivity can be enforced through the constraint $\|\mu\|_{\infty}<1$. To make this practical on the sphere, we introduce the Spherical Beltrami Differential (SBD), a two-chart representation of quasiconformal self-maps of the unit sphere $\mathbb{S}^{2}$, together with cross-chart consistency conditions that yield a globally bijective spherical deformation (up to conformal automorphisms). Building on the Spectral Beltrami Network, we develop BOOST, a differentiable optimization framework that updates two Beltrami fields to minimize task-driven losses while regularizing distortion and enforcing consistency along the seam. Experiments on large-deformation landmark matching and intensity-based spherical registration demonstrate improved task performance meanwhile maintaining controlled distortion and robust bijective behavior. We also apply the method to cortical surface registration by aligning sulcal landmarks and matching cortical sulcal depth, achieving comparative or better registration performance without sacrificing geometric validity.
@@
Quasiconformal mapping, surface registration, deep learning, surface mesh
@@ 1 Introduction
Surface registration refers to the process of establishing a meaningful bijective correspondence between two geometric surfaces. This fundamental task serves as a prerequisite for comparing shapes, transferring attributes, and analyzing geometric variations across different objects. Applications can be found in many fields, ranging from computer vision [chen2024three, li2024bi] to medical imaging and computational anatomy [chen2024multiscale, ren2024sugar]. In particular, registration between genus-0 closed surfaces has attracted significant attention in recent years due to its critical role in analyzing biological structures with spherical topology. For example, brain surface registration enables standardized comparisons between individuals for statistical analysis [lui2007landmark, ren2024sugar, zhao2022fast]. Another related application is the computation of spherical parameterizations with controlled distortion, which facilitate geometry processing and numerical analysis on genus‑0 closed surfaces [gu2004genus, choi2015flash, wang2016bijective, liao2024convergence].
The registration can be done by parameterizing the surfaces to a common template space, specifically the spherical surface. The problem of optimizing a mapping from a genus-0 closed surface to another one is then boiled down to finding an optimal sphere self-mapping. This is achieved by computing the respective spherical parameterization of each surface, whose existence is guaranteed by the uniformization theorem. Numerically, there are many established tools [choi2015flash, gu2004genus, jin2008discrete, liao2024convergence] that enable its efficient computation. Therefore, let $\phi_{i}:\mathcal{M}_{i}\to\mathbb{S}^{2}$ be the spherical parameterization of each $\mathcal{M}_{i}$, then the problem can be formulated mathematically as finding an optimal sphere self-mapping such that
@@
and the composition $\phi_{2}^{-1}\circ f\circ\phi_{1}$ is the desired mapping from $\mathcal{M}_{1}$ to $\mathcal{M}_{2}$. In this formulation, $E$ is the energy functional by which the solution satisfies desired properties, and $\mathcal{C}$ is the constraint space. In applications, bijectivity is essential for the map. Beyond bijectivity, we require the map to have well-controlled geometric distortion so that it is physically plausible. Accordingly, the feasible set $\mathcal{C}$ should be understood as bijective maps together with an explicit distortion constraint or regularization (e.g., bounds on local anisotropy), rather than arbitrary smooth bijections. For example, in the landmark constrained conformal mapping problem[lui2010optimized, lam2014landmark], to get a bijective and plausible deformation, it is necessary to control the local geometric distortion like angle distortion.
Broadly, existing approaches either optimize vertex positions directly on the sphere (e.g., energy minimization[lai2014folding] or flow-based updates[yeo2009spherical]) or pull the problem back to a planar domain via stereographic projection and optimize in $\mathbb{C}$ [choi2016spherical, choi2015flash, haker2002conformal, liao2024convergence]. While intuitive and widely used, direct optimization [lai2014folding, yeo2009spherical, zhao2022fast, zhao2019spherical] typically leads to a highly nonconvex landscape and does not inherently enforce global bijectivity. Chart-based formulations [choi2016spherical, choi2015flash, haker2002conformal, liao2024convergence, guo2023automatic] alleviate some numerical difficulty but introduce their own geometric constraints. They usually introduce artificial boundary/pole handling because the sphere cannot be covered by one chart and this can induce significant distortion near the pole. In addition, current approaches, when applied to registration problems, often struggle to simultaneously achieve both bijectivity and low geometric distortion. A more detailed review is given in Section 2.
@@
These limitations motivate our exploration of representing spherical bijection through distortion variables rather than vertex coordinates and developing optimization frameworks on the distortion variables directly so as to control the geometric distortion of the result mapping efficiently. In quasiconformal theory, the Beltrami differential/coefficient provides a compact, intrinsic description of local shear and stretch and there are mature algorithms for reconstructing quasiconformal mappings with Beltrami coefficient given [lui2012optimization, meng2016tempo, lui2013texture]. However, two key obstacles remain: (i) on the sphere, a quasiconformal representation and its computation naturally require a multi-chart viewpoint and a free-boundary mapping algorithm, and (ii) classical solvers are not designed as differentiable primitives for modern gradient-based optimization. In this work, we introduce the spherical Beltrami differential as a two-chart representation with explicit cross-chart consistency, and combine it with a differentiable quasiconformal solver surrogate, the Spectral Beltrami Network (SBN), based on least-squares quasiconformal energy (LSQC) [qiu2019computing, xu2025neural] to enable efficient, distortion-controlled optimization for spherical mapping tasks. In particular, to solve the bijective spherical parameterization problem, we propose a novel algorithm, Beltrami Optimization On Spherical Topology (BOOST). BOOST represents a spherical map using two stereographic charts (hemispheres) and optimizes the corresponding Beltrami fields using the SBN. The spherical Beltrami differential has a correspondence with the spherical quasiconformal map once cross-chart compatibility is satisfied. In practice, BOOST enforces this compatibility through seam matching, folding penalties after stitching, and seam smoothness regularization. This two-chart formulation avoids the pole distortion inherent in single-chart spherical parameterizations and enables flexible, distortion-controlled non-overlap registration on genus-0 surfaces, with applications to cortical surface alignment and landmark/feature or hybrid matching.
@@
The organization of the paper is as follows. In Section 2, related works are reviewed, and contributions are highlighted in Section 3. In Section 4, necessary mathematical concepts are introduced. In Section 5, we propose the concept of the spherical Beltrami differential and show its correspondence with spherical quasiconformal mappings, and then introduce the model Spectral Beltrami Network. In Section 6, we establish the neural-based optimization framework BOOST. In Section 7, multiple experiments are conducted to show the capability and potential of our work in complex imaging problems, with applications to cortical surface registration. In Section 8, we analyze the key factors hidden behind the excellence of the model and briefly discuss its limitations as well as potential future work. In Section 9, a conclusion is given.
@@ 2 Previous works
Spherical mapping for genus‑0 surfaces plays an important role in surface registration problems. A common pipeline is to first compute an initial spherical conformal map and then solve a spherical self-mapping problem to meet downstream goals such as distortion reduction, landmark alignment, feature matching, or image-driven registration. In this section, we review related work from three angles: (i) distortion-driven spherical parameterization, (ii) task-driven spherical registration/correspondence, and (iii) distortion control via quasiconformal/Beltrami representations.
Distortion-driven spherical parameterization
A large body of work focuses on mapping genus–0 surfaces to the sphere with bounded or optimized distortion, primarily as a parameterization or remeshing tool. Early and influential approaches include (i) direct spherical parameterization such as [praun2003spherical, wan2012efficient] based on nonlinear energies, flow based methods [jin2008discrete, kazhdan2012can], density-equalizing/area-preserving mapping [lyu2024spherical, liu2026spherical, cui2019spherical] and as-rigid-as-possible (ARAP) spherical parameterization [wang2014rigid] that trades off angle and area distortions. (ii) parameterization domain based approaches like [haker2002conformal, angenent1999conformal, angenent2002laplace] that compute conformal parameterization via solving PDE on the extended complex plane, and the subsequent improvement [choi2015flash, choi2016spherical, liao2024convergence] that propose north-south pole iterative schemes to mitigate angle distortion around poles. These methods are highly effective when the goal itself is a low-distortion parameterization, but they do not directly address general mapping tasks that must simultaneously satisfy task objectives and maintain controllable distortion.
Task-driven spherical mapping and registration
In many imaging applications, the spherical map is not the end of the story but a representation that converts surface matching to registering signals on the sphere (e.g., curvature, sulcal depth, functional features). In computational neuroanatomy, classical spherical registration is central to cortical surface analysis and atlas building. Representative pipelines include spherical coordinate systems and spherical morphing used in widely adopted tools [fischl1999high], diffeomorphic registration on the sphere such as Spherical Demons [yeo2009spherical], and flexible frameworks like Multimodal Surface Matching (MSM) that align multiple features with explicit smoothness/regularization models [robinson2014msm]. Recent years have seen significant progress in learning-based spherical surface registration. Spherical CNNs and spherical U-Net-type models introduced practical spherical convolutions for tasks such as segmentation/parcellation and spherical signal analysis [cohen2018spherical, zhao2019spherical]. Examples include SphereMorph [cheng2020cortical], S3Reg [Zhao2021S3Reg], the deep-discrete spherical registration framework (DDR) [suliman2022deep], and recent graph-based spherical registration models such as SUGAR [ren2024sugar]. These methods substantially decrease runtime and often achieve strong task accuracy. However, in these methods, the topology is preserved mainly through folding penalties and/or diffeomorphic parameterizations of the deformation model, while explicit and tunable control of geometric distortion is usually absent or only via an indirect smoothness regularization on the result mappings. This limitation becomes particularly acute when the task loss encourages highly localized stretching or twisting.
Distortion control via quasiconformal theory
Quasiconformal (QC) geometry offers a principled mechanism to quantify and constrain distortion: the Beltrami coefficient/differential encodes local shear and anisotropic stretch, and regularization on its norm can control conformality distortion. This has motivated a series of work that represents mappings in distortion variables rather than vertex coordinates, enabling more direct control of geometric validity. Examples include Teichmüller maps for landmark matching and surface registration [lui2014teichmuller], Beltrami holomorphic flow [lui2012optimization] and related Beltrami-based numerical solvers [lui2013texture, qiu2019computing, meng2016tempo], and optimization over constrained homeomorphism spaces parameterized by Beltrami coefficients [lam2014landmark, lui2015splitting]. In parallel, quasi-conformal theories’ application to spherical topology has also been studied, often through chart-based constructions [choi2015flash, choi2016spherical]. While these approaches bring explicit distortion measures into the loop, practical gaps remain when applying QC/Beltrami formulations to spherical topology: (i) the sphere cannot be covered by a single chart, so need to adopt an iterative scheme to address pole/boundary artifacts; (ii) the optimization is typically carried out with ADMM-style schemes rather than end-to-end differentiable updates, making it less convenient to integrate general task losses and to scale to modern gradient-based learning/optimization pipelines; (iii) existing QC pipelines are often formulated around explicit landmark or boundary constraints to anchor the solution, whereas many practical registration settings are landmark-free and instead rely on dense feature or intensity objectives.
Overall, prior literature provides powerful tools for (i) spherical parameterization with optimized distortion and (ii) task-driven spherical registration, but relatively few approaches can jointly optimize task objectives while enforcing bijectivity and providing explicit, effective distortion control. This gap motivates approaches that represent spherical self-maps in distortion variables (e.g., Beltrami fields), use multi-chart formulations to avoid pole artifacts, and support efficient gradient-based optimization for diverse task losses.
@@ 3 Contributions
Our work has several key contributions and novel aspects to the field of spherical bijective mapping problems, addressing the limitations of existing methods. Specifically, our contributions are as follows:
@@
    Two-chart Beltrami formulation on the sphere. We introduce the Spherical Beltrami Differential (SBD), a two-chart representation of spherical quasiconformal self-maps with cross-chart compatibility. This provides a practical parameterization of spherical self-homeomorphisms in terms of distortion variables.
@@
    BOOST: task-driven optimization of spherical self-maps in Beltrami space. We propose BOOST, a differentiable optimization framework that updates two Beltrami fields to minimize task losses while regularizing distortion and enforcing global consistency across the seam (via boundary matching, seam smoothness, and anti-folding penalties). The framework targets general spherical mapping/registration problems.
@@
    Validation on spherical registration tasks. We demonstrate the effectiveness of the proposed formulation and optimization strategy on representative tasks, including landmark matching, intensity-based spherical registration and hybrid objectives, with experiments conducted on brain cortical surface data. Results show improved task alignment with better controlled distortion and consistent bijective behavior.
@@ 4.1 Quasiconformal Mappings
(2) $\frac{\partial f}{\partial\bar{z}}=\mu(z)\frac{\partial f}{\partial z}$
for some measurable complex-valued function $\mu$ with $\|\mu\|_{\infty}<1$, where $\frac{\partial f}{\partial z}=\frac{1}{2}(f_{x}-\sqrt{-1}f_{y})$ and $\frac{\partial f}{\partial\bar{z}}=\frac{1}{2}(f_{x}+\sqrt{-1}f_{y})$. $\mu$ is called the complex dilation or Beltrami coefficient of $f$, which measures the local deviation from a conformal map. Infinitesimally, around a point $p$, $f(p+z)$ might be expressed as an affine map with negligible error $o(|z|)$:
@@
Let $f\colon(x,y)\rightarrow(u,v)$ and $\mu=\rho+\sqrt{-1}\tau$, and denote by $\mathbf{S}(2)$ the space of all $2\times 2$ symmetric positive definite matrices whose determinant is 1, then
@@
where $Q=(Q_{ij})\colon\Omega\rightarrow\mathbf{S}(2)$. Right multiplying both sides by $D_{f}(z)^{-1}$, we obtain a linear system which is the alternative formulation of the Beltrami equation:
@@
and from this we have $\mu=\frac{q_{11}-q_{22}+2\sqrt{-1}q_{12}}{q_{11}+q_{22}+2\text{sgn}(\det{D_{f}})}$. With the constraint that $Q(z)\in\mathbf{S}(2)$, we observe that $|\mu(z)|<1$ if and only if $\det D_{f}(z)>0$. By inverse function theorem, researchers would through restricting $\|\mu\|_{\infty}<1$, achieve local bijectivity and prevent folding in the triangular mesh [lui2014teichmuller, choi2015flash].